### v1

# Boosting Graph Foundation Model from Structural Perspective

Abstract
Graph foundation models have recently attracted significant attention due to its strong generalizability. Although existing methods resort to language models to learn unified semantic representations across domains, they disregard the unique structural characteristics of graphs from different domains. To address the problem, in this paper, we boost graph foundation model from structural perspective and propose BooG. The model constructs virtual super nodes to unify structural characteristics of graph data from different domains. Specifically, the super nodes fuse the information of anchor nodes and class labels, where each anchor node captures the information of a node or a graph instance to be classified. Instead of using the raw graph structure, we connect super nodes to all nodes within their neighborhood by virtual edges. This new structure allows for effective information aggregation while unifying cross-domain structural characteristics. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. Experimental results on various datasets and tasks demonstrate the superior performance of BooG. We provide our code and data here: https://anonymous.4open.science/r/BooG-EE42/.
1 Introduction
Graph learning has achieved outstanding results in a wide range of application fields in recent years, including social networks [22, 13], citation networks [18], molecular graphs [2, 33] and recommendation systems [6]. For graph data containing complex structural information, graph neural networks (GNNs) [25, 11, 29] leverage designed message-passing mechanisms to learn consistent representations for similar instances and apply them to various downstream tasks, such as node classification [11], graph classification [29] and link prediction [31].
Although general GNN models have achieved widespread success, they often face limitations due to training labels and can only perform a single task. In practical applications, accurate ground truth labels often come at a high cost [9]. To address these challenges, graph self-supervised learning (SSL) [26, 30, 9] utilizes the structure or characteristics of the data itself for learning, without the need for explicitly providing labels for the training instances. For example, graph contrastive learning (GCL) [30] supplements supervised signals by comparing the similarity and dissimilarity between different instances in the graph data. Graph SSL methods typically employ a learning pipeline involving “pre-training” and “fine-tuning”. By designing appropriate pre-training tasks, the model is enabled to generalize to downstream tasks after fine-tuning.
Although graph SSL has been demonstrated to effectively learn representations of graph data during the pre-training phase, the disparity between pre-training tasks and downstream tasks can result in suboptimal performance [16]. With the development of large language models (LLMs), prompt learning has been proposed to narrow the gap between pre-training and downstream tasks. Recently, “prompt” has been introduced into graph learning [3, 17, 23]. Such efforts aim to narrow the gap between pre-training and downstream tasks by designing unified pre-training and prompting templates. However, graph prompt learning fails to assist pre-trained models in improving domain generalization capabilities. This is because graph data from different domains often contain specific semantic characteristics. For instance, features of molecular graphs are typically vectors representing indices of nominal features of atoms, while features of citation networks often consist of bag-of-words vectors describing paper titles and abstracts.
To enhance the generalization capability of graph learning for cross-domain graph tasks, graph foundation models based on LLMs aim to learn unified semantic representations on graph data from different domains by text-attribute graphs (TAGs) [14, 28, 7]. However, these approaches are still constrained by the unique structural characteristics of graph data from different domains. For example, citation networks contain citation relationships between academic papers, often exhibiting high levels of clustering, where papers referencing each other may involve similar topics or domains. On the other hand, chemical molecular graphs typically exhibit fixed topological structures and physical-chemical properties. The structural characteristics from different domains make it challenging for models to generalize the learned structural knowledge to other domains. To further enhance the model generalizability, there naturally arises a question: Can we develop a graph foundation model that unifies cross-domain structural characteristics?
To address the aforementioned challenges, in this paper, we propose the Boosting Graph Foundation Model, namely, BooG, which achieves robust performance and generalization capabilities across various graph domains and tasks. First, BooG uses TAGs to map textual descriptions of input data to a unified space by a pre-trained language model (LM). It then introduces virtual super nodes to construct unified structural characteristics for graph data from different domains, addressing the challenge of domain-specific structural characteristics that hinders generalization. Specifically, the virtual super nodes fuse the information of both anchor nodes and class labels, where each anchor node captures the information of a node instance or a graph instance to be classified. Instead of utilizing the raw graph structure, we introduce super nodes and connect them to all the nodes within the neighborhood of an anchor by virtual edges. Details will be illustrated in Figure 3 later. This new linkage allows for effective information aggregation while disregarding the difference in structural connectivity across domains. In this way, cross-domain model generalization can be boosted from structural perspective. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. We conduct extensive experiments on six datasets with diverse domains and tasks. In most cases, our proposed approach leads the state-of-the-art competitors. We also conduct ablation studies to validate the necessity of each component in BooG.
2 Related Works
Graph Neural Network.
As an effective technique for handling graph-structured data, GNNs [25, 11, 29, 12, 4, 27] can effectively capture complex patterns based on input node features and structural information. For instance, graph attention networks (GAT) [25] utilize attention mechanisms to learn the importance of neighboring nodes and aggregate their information with learned weights. Graph Isomorphism Network (GIN) [29] adopts a unique graph encoding method that emphasizes distinguishing non-isomorphic structures, enhancing the model’s representational capacity. However, the effectiveness of GNNs often relies on a large number of training labels, implying high labeling costs. Additionally, the design of GNNs is often tailored to specific tasks, making it challenging to generalize to other downstream tasks. For instance, GNNs designed for node classification tasks often focus more on local information of the instances, while methods designed for graph classification tasks require more attention to the global information of the instances.
Graph Self-Supervised Learning.
Graph SSL [26, 30, 9] is primarily used to address the problem of lacking labels in graph learning tasks. These methods typically employ learning pipelines that include “pre-training” and “fine-tuning” stages. Through carefully designed contrastive learning or generative pre-training tasks, the model can generalize the learned knowledge to downstream tasks after fine-tuning. For example, GraphCL [30] utilizes the pre-training of graph models by applying a self-supervised contrastive learning task on learned node embeddings. DGI [26] introduces a self-supervised pre-training task aimed at maximizing the mutual information between local node views and global graph views. While graph SSL has been demonstrated to effectively learn instance representations during the pre-training phase, the disparity between pre-training tasks and downstream tasks results in suboptimal performance.
Inspired by advancements in the field of NLP, graph prompting learning [3, 17, 23] has been proposed to bridge the gap between pre-training and downstream tasks. For example, GraphPrompt [17] introduces a learnable prompt layer that can automatically identify key information in the pre-trained model for downstream tasks. However, graph prompt learning fails to assist pre-trained models in improving domain generalization capabilities.
Graph Foundation Model based on LLM.
The advent of foundation models first became a reality in NLP due to the development of LLMs. Foundation models are trained on extensive data and can adapt to a wide range of data and downstream tasks [15]. Existing graph foundation models [14, 7, 28] that integrate LLMs can generally be divided into two categories: (1) Methods with LLMs as the backbone perform various graph tasks by converting graphs into text or tokens for LLMs utilization. (2) Methods with GNN as the backbone typically use pre-trained LLMs to obtain unified representations for instances, enhancing the representation capacity of GNNs and improving generalization. For example, UniGraph [7] proposes a cascade architecture of LM and GNN, along with a self-supervised training objective based on Masked Graph Modeling (MGM), and introduces graph instruction fine-tuning using LLMs to achieve zero-shot prediction capability. OFA [14] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs. However, these approaches are still constrained by the structural characteristics of graph data from different domains.
3 Preliminaries
3.1 Text-Attributed Graphs
[…]
3.2 Learning scenarios.
In this paper, we propose a cross-domain and cross-task graph foundation model BooG. This model consists of two parts: a pre-trained encoder and a classifier implemented with an MLP. We first pre-train the encoder and freeze the pre-trained model’s parameters when generalizing to downstream tasks. To train the classifier on downstream tasks, we split the dataset into the train, validation, and test sets, denoted as $\mathcal{D}_{train}$, $\mathcal{D}_{val}$, and $\mathcal{D}_{test}$, respectively. Their label sets are denoted as $\mathcal{Y}_{train}$, $\mathcal{Y}_{val}$, and $\mathcal{Y}_{test}$. To evaluate the generalizability of BooG, we conduct experiments in three different problem settings and adopt three graph-related tasks including node classification, graph classification and link prediction. In supervised learning setting, the classifier will be trained on the training set $\mathcal{D}_{train}$, with the optimal parameters selected based on the validation set $\mathcal{D}_{val}$, and its performance will be finally evaluated on the test set $\mathcal{D}_{test}$. For few-shot learning, $N$-way $K$-shot tasks evaluate the in-context learning ability to apply the pre-learned knowledge to a new domain and task with $N$ classes, each represented by only $K$ labeled examples. For zero-shot learning, $K$ is set to 0, indicating no prior exposure to support instances from the target classes. This setting aims to evaluate a pre-trained model’s ability to generalize and apply its learned knowledge to unseen data categories.
Figure 1: The overall process of BooG. The model’s input includes text attribute graphs and class descriptions. BooG first utilizes a LM to unify different graph data and standardizes the input for node-level and graph-level tasks as sub-graphs. Subsequently, BooG introduces constructed super nodes into sub-graphs and learns to aggregate effective information within the sub-graphs. We freeze the parameters of the pre-trained model and provide the final instance representations through similarity matching. In particular, the similarity matching process can serve as zero-shot learning to predict unseen instances. For supervised learning and few-shot learning scenarios, BooG freezes the parameters of the pre-trained model and generalizes the capabilities of the pre-trained model to multiple downstream tasks by adjusting the parameters of the MLP.
4 Methodology
We propose a graph foundation model, BooG, which enhances the model generalizability across various graph domains and downstream tasks. In this section, we first introduce how to map the input TAGs from different domains to a unified feature space using a pre-trained language model. Next, we introduce the pre-training algorithm with the goal of learning expressive representations. Finally, we describe how to generalize the pre-trained models to multiple downstream graph tasks in different scenarios. The overall framework of BooG is summarized in Figure 1.
Figure 2: The text format for graph node and class node on Pubmed.
4.1 Unifying Graphs from Diverse Domains
Most GNNs have been designed for specific domains and relied heavily on pre-processed vector features as input, thereby posing challenges in terms of broader applicability and generalization. In this paper, inspired by previous works [14, 32], we introduce TAGs to describe graph nodes using raw texts, as illustrated in Figure 2. Even though the text description comes from different graph domains, we can apply a pre-trained LM to encode these texts into the same semantic space. And in BooG, we adopt Sentence Transformer [21], which has been shown to learn effective sentence embeddings. Let $t_{v}$ denote the text attributes of a graph node $v$, then it can be encoded by the pre-trained LM as follows:
$x(v)=LM(t_{v})\in\mathbb{R}^{d},$ (1)
where $x(v)$ is the output of the LM, and $d$ is the dimension of the output vector. Similarly, let $t_{l}$ be the text attributes of a class node $l$, then it can be encoded by the same pre-trained LM as follows:
$c(l)=LM(t_{l})\in\mathbb{R}^{d}.$ (2)
4.2 Unifying Structure from Diverse Domains
With the advancement of NLP, some works [14, 7] have introduced LMs into graph learning. By mapping graph data from different domains into a unified feature space through TAGs, these efforts aim to enhance the model’s generalization capabilities. However, graph data from different domains often contain unique structural information, posing challenges for existing methods to generalize across different domains. To address the challenges, we propose BooG, which introduces virtual super nodes to construct unified structural space for graph data from different domains.
Sub-graph. Considering that the input graph dataset could consist of a set of nodes for node classification and link prediction tasks, or a set of graphs for graph classification task, we unify the input graph data into sub-graphs composed of anchor nodes and their neighborhoods. Formally, we define a sub-graph $\mathcal{G}_{sub}=\left\{s,\mathcal{N}(s)\right\}$, where $s$ is an anchor node representing a node or a graph instance, and $\mathcal{N}(s)$ denotes its neighborhoods. For node-level tasks, the anchor node $s$ represents an arbitrary node in the input graph, and $\mathcal{N}(s)$ denotes the $k$-hop neighbor nodes of the anchor node $s$. The representation of the anchor node $s$ is the node features encoded by LM, i.e., $h(s)=x(s)$. For graph-level tasks, an anchor node $s$ represents a graph, so the representation of an anchor node $s$ is defined as $h(s)=R_{graph}(\mathcal{G})$, where $R_{graph}$ is a pooling function and the neighborhoods $\mathcal{N}(s)$ include all the nodes in the graph.
Figure 3: The super nodes in sub-graphs.
Super Node. We construct virtual super nodes by fusing each anchor node with all class labels, as shown in Figure 3. For each input node/graph instance, there is only one anchor node $s$ and $C$ super nodes $p$, where $C=\left|\mathcal{L}\right|$ represents the number of classes in a graph dataset. The representation of the $j$-th super node for the $i$-th node/graph can be computed as follows:
$h(p_{ij})=h(s_{i})+\alpha c(l_{j}),j\in\left[1,C\right],$ (3)
where $h(p_{ij})\in\mathbb{R}^{d}$, $h(s_{i})$ represents the representation of an anchor node $s_{i}$, $c(l_{j})$ represents the representation of the $j$-th class label and $\alpha$ is a hyper-parameter. The neighbourhood of a super node $p_{ij}$ is defined as the neighbourhood of its corresponding anchor node $s_{i}$, i.e., $\mathcal{N}(p_{ij})=\mathcal{N}(s_{i})$. We fuse class labels with anchor nodes, aiming to leverage the class labels to assist anchor nodes in identifying and aggregating relevant information within their neighborhoods. We then employ virtual super nodes to construct unified structural characteristics for graph data from different domains.
Encoder. We construct virtual edges $\bar{e}_{pu}$ between each super node $p$ and its neighbor nodes $u\in\mathcal{N}(p)$, and then learn weights for these edges with the attention mechanism. Let the weight $a_{pu}$ denote the importance of the edge $\bar{e}_{pu}$, i.e., the importance of neighbor node $u$ to super node $p$. The weights can be computed as follows:
$a_{pu}=\frac{exp(ReLU(\vec{g}^{T}\left[W_{1}h(p)\parallel W_{2}h(u)\right]))}{{\textstyle\sum_{j\in\mathcal{N}(p)}exp(ReLU(\vec{g}^{T}\left[W_{1}h(p)\parallel W_{2}h(j)\right])}},$ (4)
where $h(u)=x(u)$ represents the node features encoded by the LM, and $W_{1},W_{2}\in\mathbb{R}^{d\times d}$ and $\vec{g}\in\mathbb{R}^{2d}$ are the learnable parameters. Then, we aggregate information from neighborhoods for each super node $p$ to get the corresponding representation. The final representation of super node $p_{ij}$ can be computed as follows:
$\hat{h}(p_{ij})=W_{3}\left[\beta h(p_{ij})+(1-\beta){\textstyle\sum_{u\in\mathcal{N}(p_{ij})}a_{(p_{ij})u}h(u)}\right],$ (5)
where $W_{3}\in\mathbb{R}^{d\times d}$ is the learnable parameters and $\beta$ is a hyper-parameter. Instead of relying on the original structure, we introduce super nodes and connect them to all nodes within their neighborhood by virtual edges. This new structure allows for effective information aggregation while unifying structure space across domains, thus boosting the model generalizability.
4.3 Pre-training
To optimize the pre-trained model, we leverage a contrastive loss function $\mathcal{L}_{pre}$. The super node $p$ aggregates information from its neighbor node $u\in\mathcal{N}(u)$ based on learned weights $a_{pu}$. Intuitively, neighbors that are more similar to the super node are more important. Therefore, for each encoded super node $\hat{h}(p_{ij})$, the original super node $h(p_{ij})$ before encoding is treated as a positive sample. Meanwhile, encoded super nodes from different classes are selected as negatives. We formulate the loss as the normalized temperature-scaled cross-entropy loss [1] as follows:
$\mathcal{L}_{pre}=-\frac{1}{C}\sum_{j}^{C}\frac{1}{n}\sum_{i}^{n}log\frac{exp(sim(\hat{h}(p_{ij}),h(p_{ij})/\tau)}{\sum_{q=1,q\neq j}^{C}(exp(sim(\hat{h}(p_{ij}),\hat{h}(p_{iq}))/\tau))}$ (6)
where $C$ is the number of classes, $n$ is the number of instances, $\tau$ is a temperature hyper-parameter and sim(·) is implemented with cosine similarity.
Similarity Matching & Zero-shot learning. We employ similarity matching to select the final representation for each input graph/node. Let $z_{i}$ denote the final representation of $i$-th input graph/node, it can be computed as:
$z_{i}=\hat{h}(p_{iw}),$ (7)
where $w=\arg\max_{j}(sim(\hat{h}(p_{ij}),c(l_{j})))$, $j\in\left[1,C\right]$. Intuitively, super nodes that effectively aggregate information within their neighborhoods are most similar to their corresponding class labels. Therefore, the index $w$ based on the maximum similarity calculation is also the predicted class for node or graph classification tasks. Then, we define the prediction for the $i$-th graph/node as:
$\hat{y}_{i}=\arg\max_{j}(sim(\hat{h}(p_{ij}),c(l_{j}))),j\in\left[1,C\right].$ (8)
Since predictions based on similarity matching do not rely on labels, and the pre-trained model can generalize to unseen data. We treat the similarity matching process as zero-shot learning. It is worth noting that when evaluating the model performance, we keep the parameters of the pre-trained model frozen.
For link prediction tasks in the zero-shot setting, we compare the similarity of input node pairs. Pairs with similarity greater than a threshold $T$ are classified as having an edge; otherwise, they are considered not to have an edge. Let $\hat{e}_{ij}$ denote the link prediction result between the pair of nodes $(v_{i},v_{j})$. It can be computed as follows:
$\hat{e}_{ij}=\begin{cases}1,\text{if}~{}sim(z_{i},z_{j})>T\\ 0,\text{otherwise}\end{cases}$ (9)
4.4 Tuning for Downstream Tasks
Next, we show how we freeze the pre-trained model and adapt it to different downstream tasks on graphs. BooG uses MLP as the classifier to predict the class of instances. Formally, the final prediction for $i$-th input graph/node is given by:
$\hat{y_{i}}=\arg\max(MLP(z_{i})).$ (10)
In link prediction tasks, we standardize it as a binary classification task and employ the concatenated representation of node pairs as the input to the MLP for prediction. In both supervised and few-shot learning scenarios, we employ cross-entropy loss to train the MLP for all tasks.
4.5 Time Complexity
We analyze the time complexity of BooG in this section. For BooG, we denote the maximum sequence length of node textual feature as $L$ and the number of instances as $n$. For an input graph/node, the complexity is $O(L^{2}d+Ld^{2})$, where $d$ is the dimensionality of the embeddings. The time complexity is $O(n*(L^{2}d+Ld^{2}))$ for processing $n$ instances. The encoder’s time complexity primarily depends on the feature dimensionality and the structure of the graph. The time complexity of the attention mechanism can be expressed as $O(2d^{2}n+2dn)$. Next, the complexity of message aggregation is $O(\bar{k}dn+d^{2}n)$, where $\bar{k}$ is the average number of neighbors per node. The overall time complexity of encoder is $O(k_{1}d^{2}n+k_{2}dn)$, where $k_{1},k_{2}$ are the coefficients. For similarity matching, the time complexity is $O(dnC)$, where $C$ is the number of classes. Overall, our method BooG has a linear time complexity w.r.t. $n$.
5 Experiments
In this section, we evaluate the generalizability of BooG on both node-level and graph-level tasks.
5.1 Experimental Settings
Datasets. We conduct experiments on six benchmark datasets, which include 4 node-level graph datasets (Cora, Pubmed, ogbn-arxiv, Wiki-CS) and 2 graph-level datasets (PCBA, HIV). Due to the space limitation, we move details on datasets (Sec. A), experimental setup (Sec. B) and hyper-parameter sensitivity analysis (Sec. C) to Appendix.
Baselines. We compare BooG with six baselines, which can be categorized into four types. (1) MLP utilizes a multi-layer perceptron to extract deep features individually for each node. (2) Graph Neural Networks: GAT [25] computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them. GIN [29] enhances the representation power of GNNs by employing a distinct graph encoding method that emphasizes the discrimination of non-isomorphic structures. (3) Graph Pre-training Models: GraphCL [30] utilizes pre-training of graph models through the application of a self-discriminative contrastive learning task on learned node embeddings. DGI [26] introduces a self-supervised pre-training task that aims to maximize the mutual information between the local node view and the global graph view. (4) Graph foundation models: OFA [14] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs.
5.2 Performance Results
We next evaluate the model generalizability across various learning scenarios and downstream tasks. We employ TAGs for all methods to unify all input data. Specifically, we employ the pre-trained language models (LMs) to encode instance features as input for models. BooG and all other self-supervised learning baselines are first pre-trained on the citation network Cora. In the supervised learning setting, the parameters of pre-trained models are frozen and the MLP is trained in a supervised manner. Other baselines are trained directly on the target datasets. In the few-shot learning setting, we freeze the parameters of the pre-trained methods and train their MLPs using a $N$-way $K$-shot task. For other methods, we train them using the $N$-way $K$-shot task, i.e., they are directly trained on target datasets with labels. To accomplish this, we construct a support set by randomly selecting $K$ examples per class from the training split. For zero-shot learning, different from the above two scenarios, OFA is first pre-trained on the citation network Cora. For GraphCL and DGI, predictions are generated by comparing the similarity between the instance representations derived from their models and the class label representations encoded by the LM. Finally, we evaluate all methods on the test sets of the target datasets.
Performance analysis. Table 1, Table 2 and Table 3 report the results in supervised learning, few-shot learning, and zero-shot learning, respectively. From these tables, we observe that: (1) BooG demonstrates the best results in most cases. (2) The GNN models GAT and GIN perform better in specific domains in supervised learning and few-shot learning. For example, GAT excels on citation network datasets, while GIN is more suited to molecular graphs. However, these methods perform poorly in other domains because they are typically designed for a single graph domain, especially in few-shot learning. (3) The two self-supervised learning methods, GraphCL and DGI, exhibit poor domain generalization capabilities in few-shot learning and zero-shot learning, since they only use the raw graph structure. (4) OFA demonstrates a certain degree of domain generalization capabilities in few-shot learning, highlighting the effectiveness of in-context learning. However, BooG outperforms OFA in most cases because it constructs unified structural characteristics for graphs from different domains, thereby enhancing the model generalizability. Overall, BooG shows strong generalization ability in all scenarios, achieving results comparable to or even better than baseline methods, which proves its effectiveness.
Table 1: Experiment results in supervised learning. We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks. We highlight the best score on each dataset in bold and the runner-up score with underline.
Datasets and tasks MLP GAT GIN OFA GraphCL DGI BooG
[…]
Table 2: Experiment results in few-shot learning. We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks. We highlight the best score on each dataset in bold and the runner-up score with underline.
Datasets and tasks MLP GAT GIN OFA GraphCL DGI BooG
[…]
Table 3: Experiment results in zero-shot learning. We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks. We highlight the best score on each dataset in bold and the runner-up score with underline. Since OFA is first pre-trained on the citation network Cora with training labels, so we did not report the results of OFA on Cora in zero-shot learning.
Datasets and tasks Cora Pubmed ogbn-arxiv Wiki-CS Cora-link PCBA HIV
[…]
DGI 38.36 28.46 24.11 23.76 35.85 21.02 25.11
[…]
Figure 4: t-SNE visualization on Cora.
Figure 5: Ablation study.
5.3 Visualization
We further evaluate the expressive power of BooG. We visualize node representations before and after pre-training on Cora with the t-SNE technique [24]. Colors are used to distinguish different node classes. Before pre-training, the node representations are encoded by the LM. As shown in Figure 5, the node representations of different classes are more clearly separated after pre-training. This indicates that BooG can learn more expressive representations.
5.4 Ablation Study
The ablation study is done to understand the importance of main components of BooG in supervised learning and 5-shot learning scenarios on Cora and HIV. The model is pre-trained on Cora. We first remove the class labels from the super nodes and call this variant BooG-nc (no class label). Accordingly, in the pre-training task, we use the anchor node before aggregation as the positive sample and treat other instances of dataset as negative samples. Secondly, we remove the anchor nodes from the super nodes and treat the anchor node as a member of the contextual neighborhood. We call this variant BooG-na (no anchor node). Further, we remove the similarity matching process and concatenate the representations based on different class labels as the final representation of the input to the MLP. We call this variant BooG-ns (no similarity matching). We compare BooG with these three variants, and the results are presented in Figure 5. Our findings show that BooG outperforms all the variants on the two datasets. Further, the decline in performance exhibited by both BooG-nc and BooG-na highlights the importance of super nodes. In other words, the super nodes significantly enhance model generalizability, especially in graph domain transfer.
6 Conclusion
In this paper, we boost graph foundation model from structural perspective and propose BooG. The model constructs virtual super nodes to unify structural characteristics of graph data from different domains. Specifically, the virtual super nodes fuse the information of both anchor nodes and class labels, where each anchor node captures the information of a node instance or a graph instance to be classified. Instead of utilizing the raw graph structure, we introduce super nodes and connect them to all the nodes within the neighborhood of an anchor by virtual edges. This new linkage allows for effective information aggregation while disregarding the difference in structural connectivity across domains. In this way, cross-domain model generalization can be boosted from structural perspective. Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks. Our extensive experiments and ablation studies validate the effectiveness of BooG.
Appendix A Datasets
In this section, we introduce the datasets used in the experiments. The statistics of datasets are listed in Table 4, and we describe each dataset in detail below:
Cora is a homophilous graphs commonly used as benchmarks. It consists of 2708 scientific publications classified into one of 7 classes. In this dataset, each node represents a scientific paper and each edge represents a citation. We collect raw text from [5].
Pubmed is a homophilous graphs commonly used as benchmarks. It consists of 19,717 scientific publications from PubMed database pertaining to diabetes classified into one of 3 classes. We collect raw text from [5].
ogbn-arXiv is an academic citation network graph dataset, consisting of papers from arXiv and their citation relationships. Each node represents a paper, and edges represent the citation relationships between papers. The task is to predict the 40 subject areas of arXiv CS papers. We collect raw text from [8].
Wiki-CS is an Internet link network where each node represents a Wikipedia page and each edge represents a reference link. Each node’s label corresponds to the category of the entry. We collect raw text from [14].
PCBA is a large-scale chemical bioactivity dataset used for graph machine learning research in cheminformatics and drug discovery. Each node represents an atom of a molecule, and the edges represent the chemical bonds between atoms within the molecule. The primary tasks in the PCBA dataset are to predict the activity of chemical molecules across various bioactivity assays. We collect raw text from [14].
HIV is a subset of the PubChem BioAssay dataset, focusing on predicting the activity of chemical molecules against the HIV virus. The task in the HIV dataset involves predicting the activity of chemical molecules against the HIV virus. We collect raw text from [14].
For Cora and PubMed, we measure the performance of all models on the test sets over 10 random splits as suggested in [20] and report the average accuracy. For ogbn-arxiv, we follow the official splits [8]. Following the experimental procedure suggested by OGB, we repeat each experiment for 10 times with random seeds and report the average accuracy. For Wiki-CS, we follow the official splits [19] with 20 different training splits and report the average accuracy. For PCBA and HIV, we follow the official splits [32], repeat each experiment for 10 times with random seeds and report the average accuracy. The statistics and details of these datasets can be found in Table 4.
Table 4: Statistics of all text-attributed graph datasets.
Dataset Domain Task #Graphs Avg.#Nodes Avg.#Edges Raw Texts
Cora Citation Node/Link 1 2,708 10,556 paper titles and abstracts
Pubmed Citation Node/Link 1 19,717 44,338 paper titles and abstracts
ogbn-arxiv Citation Node 1 169,343 1,166,243 paper titles and abstracts
Wiki-CS Web link Node 1 11,701 216,123 wikipedia entry names and contents
PCBA Molecule Graph 437,929 26.0 28.1 textual descriptions of atoms/bonds
HIV Molecule Graph 41,127 25.5 27.5 textual descriptions of atoms/bonds
Appendix B Experimental Setup
We implement BooG by PyTorch and conducted the experiments with one A100 GPU. The model is optimized by Adam [10]. We employ TAGs for all methods to unify all input data. Specifically, we employ the pre-trained language models (LMs) to encode instance features as input for models. BooG and all other self-supervised learning baselines are first pre-trained on the citation network Cora. For other self-supervised learning baseline methods, we train a linear classifier on the top of the embeddings derived from the frozen model. Then, we evaluate the model performance on the test sets of all six datasets with different tasks. We perform a grid search to tune hyper-parameters based on the validation set. Details on the search space of these hyper-parameters can be found in Table 5.
Table 5: Grid search space.
Hyper-parameter Search space
lr {0.01, 0.02, 0.1, 0.2}
dropout [0.0,0.9]
weight decay {5e-5, 1e-5, 5e-4, 1e-4, 5e-3}
$\alpha$ [0.1,0.9]
$\beta$ [0.1,0.9]
$\tau$0.1,1,10}
T [0.1,0.9]
Appendix C Sensitivity Analysis
Here we study the sensitivity of two key hyper-parameters: $\alpha$ and $\beta$. The results are shown in Figure 6. From the figure, we see that the hyper-parameter $\alpha$, which represents the importance of class labels when constructing super nodes, has a significant impact on the model’s performance. If the weight of the class labels is too small, the class labels will not be effective, leading to a decline in model performance. Conversely, if the weight of the class labels is too large, the representations of the instances may lose their individuality, also resulting in a performance drop. For hyper-parameter $\beta$, BooG consistently performs well across a wide range of values, which demonstrates that $\beta$ does not significantly impact the performance of BooG.
Figure 6: Hyper-parameter sensitivity analysis.
Appendix D Broader Impacts
The primary objective of this study is to chieves robust performance and generalization capabilities across various graph domains and tasks with a graph foundation model. The results indicate that the proposed model contribute to enhance the model generalizability, which can have significant implications in various fields such as social network analysis and molecular property prediction. However, this work may also have some negative impacts. As deep neural network models become more widespread, they can be prone to misuse, potentially resulting in detrimental outcomes. Additionally, the privacy of training data remains an issue, as sensitive information may be unintentionally leaked through the model. We could incorporate responsible AI techniques for GNNs into our model to avoid the issues of misuse and privacy.
Appendix E Limitations
We discuss the limitations and future directions of our work in this section. This method is currently only suitable for graph classification, node classification and link prediction tasks. In future work, we aim to extend our method to a wider range of graph tasks. In addition, our method does not take advantage of the textual attributes of the edges, and we also aim to extend our method to more data attributes.

### v2

@@
-Boosting Graph Foundation Model from Structural Perspective
+Boosting Cross-Domain and Cross-Task Generalization for Text-Attributed Graphs from Structural Perspective
 Abstract
-Graph foundation models have recently attracted significant attention due to its strong generalizability.
-Although existing methods resort to language models to learn unified semantic representations across domains, they disregard the unique structural characteristics of graphs from different domains.
-To address the problem, in this paper, we boost graph foundation model from structural perspective and propose BooG.
+Graph models based on large language models (LLMs) have recently garnered considerable attention due to its significant success.
+Although existing methods resort to LLMs to learn unified semantic representations across domains, they disregard the unique structural characteristics of graphs from different domains.
+To address this problem, in this paper, we boost graph models from structural perspective and propose BooG.
 The model constructs virtual super nodes to unify structural characteristics of graph data from different domains.
 Specifically, the super nodes fuse the information of anchor nodes and class labels, where each anchor node captures the information of a node or a graph instance to be classified.
-Instead of using the raw graph structure, we connect super nodes to all nodes within their neighborhood by virtual edges.
-This new structure allows for effective information aggregation while unifying cross-domain structural characteristics.
+Instead of using the raw graph structure, the super nodes, along with virtual edges, establish a standardized aggregation mechanism that fuses rich information from neighborhoods and associated class labels, accommodating graph structural characteristics inherent to different domains.
 Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks.
 Experimental results on various datasets and tasks demonstrate the superior performance of BooG.
-We provide our code and data here: https://anonymous.4open.science/r/BooG-EE42/.
+We provide our code and data here: https://github.com/cy623/BooG.
+keywords:
+Graph learning, Graph foundation model
+[…]
+[…]
 1 Introduction
-Graph learning has achieved outstanding results in a wide range of application fields in recent years, including social networks [22, 13], citation networks [18], molecular graphs [2, 33] and recommendation systems [6].
-For graph data containing complex structural information, graph neural networks (GNNs) [25, 11, 29] leverage designed message-passing mechanisms to learn consistent representations for similar instances and apply them to various downstream tasks, such as node classification [11], graph classification [29] and link prediction [31].
+Graph learning has achieved outstanding results in a wide range of application fields in recent years, including social networks [1, 2], citation networks [3], molecular graphs [4, 5] and recommendation systems [6, 7].
+For graph data that encompasses intricate structural information, graph neural networks (GNNs) [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] employ sophisticated message-passing mechanisms to effectively learn consistent representations for instances that share similar characteristics, thereby enabling the application of these representations to a wide range of downstream tasks [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], such as node classification [9], graph classification [10] and link prediction [35].
 Although general GNN models have achieved widespread success, they often face limitations due to training labels and can only perform a single task.
 In practical applications, accurate ground truth labels often come at a high cost [36].
-To address these challenges, graph self-supervised learning (SSL) [26, 30, 9] utilizes the structure or characteristics of the data itself for learning, without the need for explicitly providing labels for the training instances.
+To address these challenges, graph self-supervised learning (SSL) [37, 38, 36, 39, 40, 41] utilizes the structure or characteristics of the data itself for learning, without the need for explicitly providing labels for the training instances.
 For example, graph contrastive learning (GCL) [38] supplements supervised signals by comparing the similarity and dissimilarity between different instances in the graph data.
@@
 For instance, features of molecular graphs are typically vectors representing indices of nominal features of atoms, while features of citation networks often consist of bag-of-words vectors describing paper titles and abstracts.
-To enhance the generalization capability of graph learning for cross-domain graph tasks, graph foundation models based on LLMs aim to learn unified semantic representations on graph data from different domains by text-attribute graphs (TAGs) [14, 28, 7].
+To enhance the generalization capability of graph learning for cross-domain graph tasks, graph models based on LLMs aim to learn unified semantic representations on graph data from different domains by text-attribute graphs (TAGs) [46, 47, 48].
 However, these approaches are still constrained by the unique structural characteristics of graph data from different domains.
@@
 The structural characteristics from different domains make it challenging for models to generalize the learned structural knowledge to other domains.
-To further enhance the model generalizability, there naturally arises a question: Can we develop a graph foundation model that unifies cross-domain structural characteristics?
-To address the aforementioned challenges, in this paper, we propose the Boosting Graph Foundation Model, namely, BooG, which achieves robust performance and generalization capabilities across various graph domains and tasks.
-First, BooG uses TAGs to map textual descriptions of input data to a unified space by a pre-trained language model (LM).
-It then introduces virtual super nodes to construct unified structural characteristics for graph data from different domains, addressing the challenge of domain-specific structural characteristics that hinders generalization.
-Specifically, the virtual super nodes fuse the information of both anchor nodes and class labels, where each anchor node captures the information of a node instance or a graph instance to be classified.
-Instead of utilizing the raw graph structure, we introduce super nodes and connect them to all the nodes within the neighborhood of an anchor by virtual edges.
-Details will be illustrated in Figure 3 later.
-This new linkage allows for effective information aggregation while disregarding the difference in structural connectivity across domains.
-In this way, cross-domain model generalization can be boosted from structural perspective.
-Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks.
-We conduct extensive experiments on six datasets with diverse domains and tasks.
-In most cases, our proposed approach leads the state-of-the-art competitors.
+To address the aforementioned challenges, in this paper, we propose BooG, a pre-trained graph model designed for Boosting cross-domain and cross-task generalization for text-attributed Graphs.
+BooG unifies the attributes and structures of graph data across different domains, thereby enhancing the capabilities of the pre-trained model to downstream tasks and data from various domains.
+Specifically, BooG first leverages a pre-trained language model (LM) to encode the attributes of the graph data and class labels into a unified representation space.
+Next, BooG unifies graph structures by introducing sub-graphs and super nodes.
+Sub-graphs offer a consistent task template across all levels of tasks, consisting of anchor nodes and their neighbor nodes.
+The super nodes, along with virtual edges, establish a standardized aggregation mechanism that fuses rich information from neighborhoods and associated class labels, accommodating graph structural characteristics inherent to different domains.
+Additionally, we propose a novel pre-training objective based on contrastive learning, which enables the model to learn more expressive nodes/graphs representations that generalize effectively to different downstream domains and tasks.
+Our extensive experiments on seven datasets, covering a range of domains and tasks, show that BooG outperforms state-of-the-art competitors in most cases.
 We also conduct ablation studies to validate the necessity of each component in BooG.
 2 Related Works
-Graph Neural Network.
-As an effective technique for handling graph-structured data, GNNs [25, 11, 29, 12, 4, 27] can effectively capture complex patterns based on input node features and structural information.
+2.1 Graph Neural Network.
+As an effective technique for handling graph-structured data, GNNs [8, 9, 10, 49, 50, 51, 52, 53, 54, 55, 56, 26, 57] can effectively capture complex patterns based on input node features and structural information.
 For instance, graph attention networks (GAT) [8] utilize attention mechanisms to learn the importance of neighboring nodes and aggregate their information with learned weights.
@@
 For instance, GNNs designed for node classification tasks often focus more on local information of the instances, while methods designed for graph classification tasks require more attention to the global information of the instances.
-Graph Self-Supervised Learning.
-Graph SSL [26, 30, 9] is primarily used to address the problem of lacking labels in graph learning tasks.
+2.2 Graph Self-Supervised Learning.
+Graph SSL [37, 38, 36, 39, 40, 41] is primarily used to address the problem of lacking labels in graph learning tasks.
 These methods typically employ learning pipelines that include “pre-training” and “fine-tuning” stages.
@@
 However, graph prompt learning fails to assist pre-trained models in improving domain generalization capabilities.
-Graph Foundation Model based on LLM.
+2.3 Graph Foundation Model based on LLM.
 The advent of foundation models first became a reality in NLP due to the development of LLMs.
 Foundation models are trained on extensive data and can adapt to a wide range of data and downstream tasks [58].
-Existing graph foundation models [14, 7, 28] that integrate LLMs can generally be divided into two categories: (1) Methods with LLMs as the backbone perform various graph tasks by converting graphs into text or tokens for LLMs utilization.
+Existing graph models [46, 48, 47, 59, 60, 61] that integrate LLMs can generally be divided into two categories: (1) Methods with LLMs as the backbone perform various graph tasks by converting graphs into text or tokens for LLMs utilization.
 (2) Methods with GNN as the backbone typically use pre-trained LLMs to obtain unified representations for instances, enhancing the representation capacity of GNNs and improving generalization.
@@
 OFA [46] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs.
-However, these approaches are still constrained by the structural characteristics of graph data from different domains.
+However, these approaches are still constrained by structural characteristics of graph from different domains.
 3 Preliminaries
@@
 […]
-3.2 Learning scenarios.
-In this paper, we propose a cross-domain and cross-task graph foundation model BooG.
-This model consists of two parts: a pre-trained encoder and a classifier implemented with an MLP.
-We first pre-train the encoder and freeze the pre-trained model’s parameters when generalizing to downstream tasks.
-To train the classifier on downstream tasks, we split the dataset into the train, validation, and test sets, denoted as $\mathcal{D}_{train}$, $\mathcal{D}_{val}$, and $\mathcal{D}_{test}$, respectively.
+Figure 1: The overall process of BooG.
+The model consists of two parts: A pre-trained model based on self-supervised manner and a downstream classifier implemented with a MLP.
+The model’s input includes text attribute graphs and class descriptions.
+BooG first utilizes a pre-trained LM to unify different graph data and standardizes the input for node-level and graph-level tasks as sub-graphs.
+Subsequently, BooG introduces super nodes to establish a standardized aggregation mechanism that fuses rich information from neighborhoods and associated class labels.
+We freeze the parameters of the pre-trained model and provide the final instance representations through similarity matching.
+In particular, the similarity matching process can serve as zero-shot learning to predict unseen instances.
+For supervised learning and few-shot learning scenarios, BooG freezes the parameters of the pre-trained model and generalizes the capabilities of the pre-trained model to multiple downstream tasks by adjusting the parameters of the MLP.
+3.2 Message Propagation
+Most graph neural networks follow the message-passing mechanism, which consists of two main steps: aggregation and update.
+We take an arbitrary node $v_{i}$ in the $k$-th layer as an example.
+The first sub-step is to aggregate information from a node’s neighbors, which is given as:
+$\hat{h}_{i}^{(k)}=\texttt{AGGREGATE}(h_{j}^{(k-1)},\forall v_{j}\in\mathcal{N}_{i}).$ (1)
+After that, the second sub-step is to update node embeddings:
+$h_{i}^{(k)}=\texttt{UPDATE}(h_{i}^{(k-1)},\hat{h}_{i}^{(k)}).$ (2)
+In this paper, we implement the encoder by learning information aggregation on the newly constructed structure.
+3.3 Learning Scenarios
+In this paper, we propose a cross-domain and cross-task graph model BooG.
+This model consists of two parts: A pre-trained model based on self-supervised manner and a downstream classifier implemented with a Multi-Layer Perceptron (MLP).
+The pre-trained model, denoted as $f_{\theta}$, operates on a TAG and generates an embedding for each sample (a node or a graph).
+We freeze the pre-trained model’s parameters when generalizing to downstream tasks.
+Our model’s generalization ability is evaluated across domains using unseen datasets through three different machine learning problems, and we adopt three graph-related tasks, including node classification, graph classification, and link prediction.
+We split the dataset into the train, validation, and test sets, denoted as $\mathcal{D}_{train}$, $\mathcal{D}_{val}$, and $\mathcal{D}_{test}$, respectively.
 Their label sets are denoted as $\mathcal{Y}_{train}$, $\mathcal{Y}_{val}$, and $\mathcal{Y}_{test}$.
-To evaluate the generalizability of BooG, we conduct experiments in three different problem settings and adopt three graph-related tasks including node classification, graph classification and link prediction.
 In supervised learning setting, the classifier will be trained on the training set $\mathcal{D}_{train}$, with the optimal parameters selected based on the validation set $\mathcal{D}_{val}$, and its performance will be finally evaluated on the test set $\mathcal{D}_{test}$.
@@
 This setting aims to evaluate a pre-trained model’s ability to generalize and apply its learned knowledge to unseen data categories.
-Figure 1: The overall process of BooG.
-The model’s input includes text attribute graphs and class descriptions.
-BooG first utilizes a LM to unify different graph data and standardizes the input for node-level and graph-level tasks as sub-graphs.
-Subsequently, BooG introduces constructed super nodes into sub-graphs and learns to aggregate effective information within the sub-graphs.
-We freeze the parameters of the pre-trained model and provide the final instance representations through similarity matching.
-In particular, the similarity matching process can serve as zero-shot learning to predict unseen instances.
-For supervised learning and few-shot learning scenarios, BooG freezes the parameters of the pre-trained model and generalizes the capabilities of the pre-trained model to multiple downstream tasks by adjusting the parameters of the MLP.
 4 Methodology
-We propose a graph foundation model, BooG, which enhances the model generalizability across various graph domains and downstream tasks.
-In this section, we first introduce how to map the input TAGs from different domains to a unified feature space using a pre-trained language model.
-Next, we introduce the pre-training algorithm with the goal of learning expressive representations.
-Finally, we describe how to generalize the pre-trained models to multiple downstream graph tasks in different scenarios.
-The overall framework of BooG is summarized in Figure 1.
-Figure 2: The text format for graph node and class node on Pubmed.
-4.1 Unifying Graphs from Diverse Domains
-Most GNNs have been designed for specific domains and relied heavily on pre-processed vector features as input, thereby posing challenges in terms of broader applicability and generalization.
-In this paper, inspired by previous works [14, 32], we introduce TAGs to describe graph nodes using raw texts, as illustrated in Figure 2.
-Even though the text description comes from different graph domains, we can apply a pre-trained LM to encode these texts into the same semantic space.
-And in BooG, we adopt Sentence Transformer [21], which has been shown to learn effective sentence embeddings.
+In this section, we present the BooG framework, as illustrated in Figure 1.
+Figure 2: The text format for graph node and class label on Pubmed.
+4.1 Unifying Graph Attributes
+Most learned GNNs are tailored to specific domains and rely on pre-processed vector features, which limits their ability to generalize across domains due to the inherently distinct and domain-specific nature of training graph features.
+Fortunately, since graph attributes can be expressed in natural language [46], in this paper, we unify graph data attributes from different domains by converting input graphs into TAGs and encoding their semantic information using a pre-trained LM.
+This transformation enables graph nodes to be represented as raw text, as illustrated in Figure 2.
+Specifically, we employ the Sentence Transformer [62] as our pre-trained LM, known for its capability to generate effective sentence embeddings.
 Let $t_{v}$ denote the text attributes of a graph node $v$, then it can be encoded by the pre-trained LM as follows:
-$x(v)=LM(t_{v})\in\mathbb{R}^{d},$ (1)
+$x(v)=\text{LM}(t_{v})\in\mathbb{R}^{d},$ (3)
 where $x(v)$ is the output of the LM, and $d$ is the dimension of the output vector.
-Similarly, let $t_{l}$ be the text attributes of a class node $l$, then it can be encoded by the same pre-trained LM as follows:
-$c(l)=LM(t_{l})\in\mathbb{R}^{d}.$ (2)
-4.2 Unifying Structure from Diverse Domains
-With the advancement of NLP, some works [14, 7] have introduced LMs into graph learning.
-By mapping graph data from different domains into a unified feature space through TAGs, these efforts aim to enhance the model’s generalization capabilities.
-However, graph data from different domains often contain unique structural information, posing challenges for existing methods to generalize across different domains.
-To address the challenges, we propose BooG, which introduces virtual super nodes to construct unified structural space for graph data from different domains.
+In addition, each class $l\in\mathcal{L}$ of each dataset is associated with a text attribute $t_{l}\in\mathcal{T_{L}}$.
+The text attributes $t_{l}$ are also encoded by the same pre-trained LM as follows:
+$c(l)=\text{LM}(t_{l})\in\mathbb{R}^{d}.$ (4)
+It is worth noting that during BooG’s pre-training phase, we rely solely on the textual descriptions of the dataset classes, rather than the class labels of nodes or graphs, thereby establishing a self-supervised learning paradigm, which will be detailed in the following sections.
+4.2 Unifying Graph Structures
+Graph datasets from different domains exhibit diverse structural characteristics.
+For example, citation networks often show high clustering coefficients, whereas molecular graphs have fixed, small-world topologies.
+Moreover, downstream tasks on these datasets can be at different granularities, such as node classification, link prediction, or graph classification.
+These discrepancies hinder a unified representation learning framework.
 Sub-graph.
-Considering that the input graph dataset could consist of a set of nodes for node classification and link prediction tasks, or a set of graphs for graph classification task, we unify the input graph data into sub-graphs composed of anchor nodes and their neighborhoods.
-Formally, we define a sub-graph $\mathcal{G}_{sub}=\left\{s,\mathcal{N}(s)\right\}$, where $s$ is an anchor node representing a node or a graph instance, and $\mathcal{N}(s)$ denotes its neighborhoods.
-For node-level tasks, the anchor node $s$ represents an arbitrary node in the input graph, and $\mathcal{N}(s)$ denotes the $k$-hop neighbor nodes of the anchor node $s$.
-The representation of the anchor node $s$ is the node features encoded by LM, i.e., $h(s)=x(s)$.
-For graph-level tasks, an anchor node $s$ represents a graph, so the representation of an anchor node $s$ is defined as $h(s)=R_{graph}(\mathcal{G})$, where $R_{graph}$ is a pooling function and the neighborhoods $\mathcal{N}(s)$ include all the nodes in the graph.
-Figure 3: The super nodes in sub-graphs.
-Super Node.
-We construct virtual super nodes by fusing each anchor node with all class labels, as shown in Figure 3.
-For each input node/graph instance, there is only one anchor node $s$ and $C$ super nodes $p$, where $C=\left|\mathcal{L}\right|$ represents the number of classes in a graph dataset.
-The representation of the $j$-th super node for the $i$-th node/graph can be computed as follows:
-$h(p_{ij})=h(s_{i})+\alpha c(l_{j}),j\in\left[1,C\right],$ (3)
-where $h(p_{ij})\in\mathbb{R}^{d}$, $h(s_{i})$ represents the representation of an anchor node $s_{i}$, $c(l_{j})$ represents the representation of the $j$-th class label and $\alpha$ is a hyper-parameter.
-The neighbourhood of a super node $p_{ij}$ is defined as the neighbourhood of its corresponding anchor node $s_{i}$, i.e., $\mathcal{N}(p_{ij})=\mathcal{N}(s_{i})$.
-We fuse class labels with anchor nodes, aiming to leverage the class labels to assist anchor nodes in identifying and aggregating relevant information within their neighborhoods.
-We then employ virtual super nodes to construct unified structural characteristics for graph data from different domains.
+To accommodate varied domains and tasks under a unified framework, we first convert each input graph into a sub-graph centering on an anchor node.
+Formally, for a given anchor node $s$, we construct a sub-graph $G_{\text{sub}}=\{s,\mathcal{N}(s)\}$, where $\mathcal{N}(s)$ denotes the neighborhood nodes of $s$.
+  • For node/edge-level tasks, $s$ is a node from the input graph; $\mathcal{N}(s)$ comprises its $k$-hop neighbors.
+  • For graph-level tasks, $s$ represents the whole graph through a pooled embedding $h(s)=R_{\text{graph}}(G)$, and $\mathcal{N}(s)$ contains all nodes in the graph.
+This step ensures that all tasks are reduced to a common format centered around an anchor node and its context.
+Figure 3: Unifying graph structures by super nodes.
+Super Nodes. We propose super nodes to explicitly incorporate class label information into the graph structure and to align structure learning across different domains and tasks.
+For each anchor node $s_{i}$ in a sub-graph, we create $C$ super nodes, where $C=|\mathcal{L}|$ is the number of class labels in the dataset.
+The $j$-th super node $p_{ij}$ is a virtual node representing the anchor node $s_{i}$ fused with the class label $l_{j}$.
+The representation is computed as:
+$h(p_{ij})=h(s_{i})+\alpha\cdot c(l_{j}),$ (5)
+where $h(s_{i})\in\mathbb{R}^{d}$ is the embedding of anchor node $s_{i}$ obtained via the language model, $c(l_{j})\in\mathbb{R}^{d}$ is the embedding of the $j$-th class label, and $\alpha$ is a hyperparameter controlling the weight of class information.
+The neighborhood of $p_{ij}$ is defined to be the same as that of $s_{i}$, i.e., $\mathcal{N}(p_{ij})=\mathcal{N}(s_{i})$.
+This design allows the super node to aggregate from the same context as the anchor node, while being explicitly tied to a specific class label.
+Anchor Node $s_{i}$ captures instance-specific information.
+Super Node $p_{ij}$ hypothesizes “what if $s_{i}$ belongs to class $l_{j}$”.
+Multiple Super Nodes provide multiple hypotheses per anchor node, one for each possible class.
+These super nodes act as explicit class-guided targets during contrastive learning.
+Aggregating neighborhood information conditioned on class embeddings facilitates more discriminative and generalizable representations.
+Unified Structure with Virtual Edges.
+The introduction of super nodes enables a standardized, super node-centered structure that unifies the aggregation process across tasks and domains.
+For each anchor node $s$, we insert all of its corresponding super nodes into its neighborhood, as illustrated in Figure 3.
+The neighborhood of each anchor node is obtained through sub-graph sampling to preserve its original structural information.
+Super nodes combine class labels with anchor nodes, leveraging these labels to assist anchor nodes in identifying relevant information within their neighborhoods.
+To explicitly construct this unified structure, we introduce virtual edges $\bar{e}_{pu}$ connecting each super node $p$ to its neighborhood nodes $u\in\mathcal{N}(p)$.
+These virtual edges differ fundamentally from the original edges in the graph $G$ in both purpose and construction.
+Original edges reflect the inherent data topology (e.g., citations or chemical bonds), whereas virtual edges are artificially constructed to support standardized aggregation conditioned on class labels.
+Specifically, for each super node $p$, $\mathcal{N}(p)$ inherits from the neighborhood of its anchor node $s$, ensuring super nodes access the same neighborhood information.
+Original edges propagate node-level information through message passing based on fixed topology.
+In contrast, virtual edges allow super nodes to selectively aggregate neighborhood information through learnable attention weights, guiding the prediction towards the correct class label.
+This design ensures that the final representations of super nodes are both instance-aware and class-informed.
+The virtual edges play a crucial role in BooG’s unified structure design, enabling consistent and comparable aggregation across different datasets and tasks.
+While original edges maintain the graph’s inherent topology, virtual edges provide a controlled mechanism for super nodes to gather label-relevant neighborhood information, which is essential for BooG’s generalization capability.
 Encoder.
-We construct virtual edges $\bar{e}_{pu}$ between each super node $p$ and its neighbor nodes $u\in\mathcal{N}(p)$, and then learn weights for these edges with the attention mechanism.
-Let the weight $a_{pu}$ denote the importance of the edge $\bar{e}_{pu}$, i.e., the importance of neighbor node $u$ to super node $p$.
-The weights can be computed as follows:
-$a_{pu}=\frac{exp(ReLU(\vec{g}^{T}\left[W_{1}h(p)\parallel W_{2}h(u)\right]))}{{\textstyle\sum_{j\in\mathcal{N}(p)}exp(ReLU(\vec{g}^{T}\left[W_{1}h(p)\parallel W_{2}h(j)\right])}},$ (4)
-where $h(u)=x(u)$ represents the node features encoded by the LM, and $W_{1},W_{2}\in\mathbb{R}^{d\times d}$ and $\vec{g}\in\mathbb{R}^{2d}$ are the learnable parameters.
-Then, we aggregate information from neighborhoods for each super node $p$ to get the corresponding representation.
-The final representation of super node $p_{ij}$ can be computed as follows:
+We then learn to aggregate information within the newly constructed structure using an attention mechanism.
+Let $a_{pu}$ denotes the weight of the virtual edge $\bar{e}_{pu}$, indicating the importance of neighbor node $u$ to super node $p$.
+The weights are calculated as follows:
+$a_{pu}=\frac{\exp\left(\text{ReLU}(\vec{g}^{\,T}\left[W_{1}h(p)\parallel W_{2}h(u)\right])\right)}{{\textstyle\sum_{j\in\mathcal{N}(p)}\exp\left(\text{ReLU}(\vec{g}^{\,T}\left[W_{1}h(p)\parallel W_{2}h(j)\right]\right)}},$ (6)
+where $h(u)=x(u)$ represents the node features encoded by the LM, $W_{1},W_{2}\in\mathbb{R}^{d\times d}$ and $\vec{g}\in\mathbb{R}^{2d}$ are learnable parameters, and $\parallel$ denotes concatenation.
+Next, we aggregate information for each super node $p_{ij}$ from its neighborhood to obtain its final representation $\hat{h}(p_{ij})$, computed as:
 $\hat{h}(p_{ij})=W_{3}\left[\beta h(p_{ij})+(1-\beta){\textstyle\sum_{u\in\mathcal{N}(p_{ij})}a_{(p_{ij})u}h(u)}\right],$ (7)
-where $W_{3}\in\mathbb{R}^{d\times d}$ is the learnable parameters and $\beta$ is a hyper-parameter.
-Instead of relying on the original structure, we introduce super nodes and connect them to all nodes within their neighborhood by virtual edges.
-This new structure allows for effective information aggregation while unifying structure space across domains, thus boosting the model generalizability.
-4.3 Pre-training
-To optimize the pre-trained model, we leverage a contrastive loss function $\mathcal{L}_{pre}$.
-The super node $p$ aggregates information from its neighbor node $u\in\mathcal{N}(u)$ based on learned weights $a_{pu}$.
-Intuitively, neighbors that are more similar to the super node are more important.
-Therefore, for each encoded super node $\hat{h}(p_{ij})$, the original super node $h(p_{ij})$ before encoding is treated as a positive sample.
-Meanwhile, encoded super nodes from different classes are selected as negatives.
-We formulate the loss as the normalized temperature-scaled cross-entropy loss [1] as follows:
-$\mathcal{L}_{pre}=-\frac{1}{C}\sum_{j}^{C}\frac{1}{n}\sum_{i}^{n}log\frac{exp(sim(\hat{h}(p_{ij}),h(p_{ij})/\tau)}{\sum_{q=1,q\neq j}^{C}(exp(sim(\hat{h}(p_{ij}),\hat{h}(p_{iq}))/\tau))}$ (6)
-where $C$ is the number of classes, $n$ is the number of instances, $\tau$ is a temperature hyper-parameter and sim(·) is implemented with cosine similarity.
-Similarity Matching & Zero-shot learning.
-We employ similarity matching to select the final representation for each input graph/node.
-Let $z_{i}$ denote the final representation of $i$-th input graph/node, it can be computed as:
-$z_{i}=\hat{h}(p_{iw}),$ (7)
-where $w=\arg\max_{j}(sim(\hat{h}(p_{ij}),c(l_{j})))$, $j\in\left[1,C\right]$.
-Intuitively, super nodes that effectively aggregate information within their neighborhoods are most similar to their corresponding class labels.
-Therefore, the index $w$ based on the maximum similarity calculation is also the predicted class for node or graph classification tasks.
-Then, we define the prediction for the $i$-th graph/node as:
-$\hat{y}_{i}=\arg\max_{j}(sim(\hat{h}(p_{ij}),c(l_{j}))),j\in\left[1,C\right].$ (8)
-Since predictions based on similarity matching do not rely on labels, and the pre-trained model can generalize to unseen data.
-We treat the similarity matching process as zero-shot learning.
-It is worth noting that when evaluating the model performance, we keep the parameters of the pre-trained model frozen.
+where $W_{3}\in\mathbb{R}^{d\times d}$ is a learnable parameters and $\beta$ is a hyper-parameter.
+4.3 Self-Supervised Learning Paradigm
+Pre-training. We pre-train BooGvia a contrastive learning objective to enhance its ability to distinguish between different class hypotheses and improve representation generalizability.
+For each anchor node $s_{i}$, we generate $C$ super nodes $\{p_{ij}\}_{j=1}^{C}$ corresponding to all possible class labels.
+These super nodes incorporate both the anchor node’s information and individual class semantics through Equation (5).
+After neighborhood aggregation, we obtain the updated embedding $\hat{h}(p_{ij})$ for each super node $p_{ij}$.
+To guide the model toward learning class-discriminative representations, we adopt a normalized temperature-scaled cross-entropy loss [63], formulated as:
+$\begin{split}&\mathcal{L}_{pre}=\\ &-\frac{1}{C}\sum_{j}^{C}\frac{1}{n}\sum_{i}^{n}log\frac{\exp(\text{sim}(\hat{h}(p_{ij}),h(p_{ij}))/\tau)}{\sum_{q=1,q\neq j}^{C}(\exp(\text{sim}(\hat{h}(p_{ij}),\hat{h}(p_{iq}))/\tau))}.\end{split}$ (8)
+Here, $n$ denotes the number of training instances, $C$ is the number of classes, $\tau$ is a temperature hyper-parameter, and $\text{sim}(\cdot)$ computes cosine similarity.
+Notably, while the loss function considers the number of classes $C$ in the dataset, it does not necessitate concrete sample-label pairs.
+Our pre-training strategy leverages self-supervised learning via contrastive learning.
+In this framework, the original embedding $h(p_{ij})$ of super node $p_{ij}$—formed by the combination of the anchor node $s_{i}$ and class label $l_{j}$—is treated as the positive sample for its updated embedding $\hat{h}(p_{ij})$.
+Meanwhile, embeddings $\hat{h}(p_{iq})$ with $q\neq j$ serve as negative samples for the same anchor node $s_{i}$, as they are associated with different class labels.
+This contrastive objective explicitly drives the model to align the aggregated representation $\hat{h}(p_{ij})$ closer to its corresponding class label (positive pair) while pushing it away from other class hypotheses (negative pairs).
+This alignment forces the model to learn more discriminative features that capture class semantics, thereby enhancing its expressive power and its ability to generalize across tasks and domains.
+Similarity Matching.
+After pre-training, each anchor node $s_{i}$ produces $C$ super node embeddings, one for each class label.
+To select the most appropriate representation for downstream tasks, we apply similarity matching:
+$z_{i}=\hat{h}(p_{iw}),\quad\text{where}\quad w=\arg\max_{j\in[1,C]}\text{sim}(\hat{h}(p_{ij}),c(l_{j})),$ (9)
+where $c(l_{j})$ is the embedding of the $j$-th class label.
+This step ensures that the final embedding $z_{i}$ for $s_{i}$ is selected based on the class label most semantically aligned with the aggregated neighborhood information.
+4.4 Tuning for Downstream Tasks
+Zero-shot Scenarios.
+Since the predictions provided by similarity matching do not depend on labels, the pre-trained model can generalize to unseen data, allowing us to treat this process as zero-shot learning.
+It is worth noting that the parameters of the pre-trained model are kept frozen during evaluation under all learning scenarios.
+For node/graph classification tasks, the prediction for the $i$-th node/graph is defined as:
+$\hat{y}_{i}=\arg\max_{j}(\text{sim}(\hat{h}(p_{ij}),c(l_{j}))),j\in\left[1,C\right].$ (10)
 For link prediction tasks in the zero-shot setting, we compare the similarity of input node pairs.
-Pairs with similarity greater than a threshold $T$ are classified as having an edge; otherwise, they are considered not to have an edge.
-Let $\hat{e}_{ij}$ denote the link prediction result between the pair of nodes $(v_{i},v_{j})$.
-It can be computed as follows:
-$\hat{e}_{ij}=\begin{cases}1,\text{if}~{}sim(z_{i},z_{j})>T\\ 0,\text{otherwise}\end{cases}$ (9)
-4.4 Tuning for Downstream Tasks
-Next, we show how we freeze the pre-trained model and adapt it to different downstream tasks on graphs.
-BooG uses MLP as the classifier to predict the class of instances.
-Formally, the final prediction for $i$-th input graph/node is given by:
-$\hat{y_{i}}=\arg\max(MLP(z_{i})).$ (10)
-In link prediction tasks, we standardize it as a binary classification task and employ the concatenated representation of node pairs as the input to the MLP for prediction.
-In both supervised and few-shot learning scenarios, we employ cross-entropy loss to train the MLP for all tasks.
+Pairs with a similarity exceeding a threshold $T$ are classified as having an edge; otherwise, they are considered not to have an edge.
+The link prediction result $\hat{e}_{ij}$ between the pair of nodes $(v_{i},v_{j})$ is computed as follows:
+$\hat{e}_{ij}=\begin{cases}1,\text{if}~\text{sim}(z_{i},z_{j})>T\\ 0,\text{otherwise}\end{cases}.$ (11)
+Supervised & Few-shot Scenarios. BooG uses MLP as the classifier to predict the class of instances. Formally, the final prediction for $i$-th input node/graph is given by:
+$\hat{y_{i}}=\arg\max(\text{MLP}(z_{i})).$ (12)
+For link prediction tasks, the problem is framed as a binary classification task, where the concatenated representation of node pairs is used as input to the MLP for prediction.
+In both supervised and few-shot learning scenarios, we employ cross-entropy loss to train the MLP across all tasks.
 4.5 Time Complexity
@@
 For similarity matching, the time complexity is $O(dnC)$, where $C$ is the number of classes.
-Overall, our method BooG has a linear time complexity w.r.t.
+Overall, BooG has a linear time complexity w.r.t.
 $n$.
+Table 1: Statistics of all text-attributed graph datasets.
+Dataset Domain Task #Graphs Avg.#Nodes Avg.#Edges Raw Texts
+Cora Citation Node/Link 1 2,708 10,556 paper titles and abstracts
+Pubmed Citation Node/Link 1 19,717 44,338 paper titles and abstracts
+ogbn-arxiv Citation Node 1 169,343 1,166,243 paper titles and abstracts
+Wiki-CS Web link Node 1 11,701 216,123 wikipedia entry names and contents
+Ele-Fashion Product Node 1 97,766 199,602 fashion titles
+PCBA Molecule Graph 437,929 26.0 28.1 textual descriptions of atoms/bonds
+HIV Molecule Graph 41,127 25.5 27.5 textual descriptions of atoms/bonds
+4.6 Why BooG Works?
+BooG boosts model’s cross-domain and cross-task generalization ability from two key aspects.
+First, BooG unifies graph attributes by encoding node texts of TAGs and class label descriptions into a shared space using a pre-trained LM, thereby equipping the model to handle graph data from diverse domains.
+More critically, BooG unifies graph structures by introducing the concept of sub-graphs and super nodes.
+Sub-graphs offer a consistent task template across all levels of tasks, consisting of anchor nodes and their neighbor nodes.
+The super nodes, along with virtual edges, establish a standardized aggregation mechanism that fuses rich information from neighborhoods and associated class labels, accommodating graph structural characteristics inherent to different domains.
+Consequently, these designs integrate a supernode-centered graph structure that encapsulates label preferences, making BooG compatible with different graph domains and task levels.
+Recent studies on graph domain generalization include UniGraph [48] and OFA [46].
+UniGraph proposes a cascade architecture integrating LM and GNN, utilizing a self-supervised objective based on Masked Graph Modeling.
+Additionally, it incorporates graph instruction fine-tuning through LLMs to enable zero-shot prediction.
+BooG freezes LM throughout the process and fundamentally resorts to the unification of graph attributes and structures as the primary vehicle for generalization.
+In addition, while OFA leverages an LM to encode node features across domains and introduces the nodes-of-interest (NOI) subgraph and NOI prompt node to unify different graph tasks, BooG’s super nodes are designed to bind anchor nodes directly with class labels, as expressed in Eq.(5), which aligns them explicitly within the same latent space, ensuring a more cohesive integration between node features and task-specific labels.
 5 Experiments
-In this section, we evaluate the generalizability of BooG on both node-level and graph-level tasks.
 5.1 Experimental Settings
-Datasets.
-We conduct experiments on six benchmark datasets, which include 4 node-level graph datasets (Cora, Pubmed, ogbn-arxiv, Wiki-CS) and 2 graph-level datasets (PCBA, HIV).
-Due to the space limitation, we move details on datasets (Sec.
-A), experimental setup (Sec.
-B) and hyper-parameter sensitivity analysis (Sec.
-C) to Appendix.
+Datasets. We first introduce the datasets used in the experiments. The statistics of these datasets are summarized in Table 1, and details of each dataset are given below:
+Cora is a classic graph dataset widely used in research on GNNs and node classification tasks.
+It consists of 2708 scientific publications classified into one of 7 classes.
+In this dataset, each node represents a scientific paper and each edge represents a citation.
+The label of nodes indicates the research field of the paper.
+We collect raw text from [64].
+Pubmed is another widely used dataset for graph-based machine learning, particularly in the context of node classification tasks with GNNs.
+It consists of 19,717 scientific publications classified into one of 3 classes.
+Similar to the Cora dataset, it is a citation network of scientific papers, but it consists of biomedical papers indexed in PubMed, which is a free search engine accessing primarily life sciences and biomedical literature.
+We collect raw text from [64].
+ogbn-arXiv is part of the Open Graph Benchmark (OGB), a collection of large-scale graph datasets designed to evaluate machine learning algorithms, particularly for graph-based learning tasks.
+It is focused on node classification in a citation network of academic papers, specifically from the arXiv repository, which is a preprint server for scientific papers in fields such as physics, mathematics, computer science, and more.
+Each node represents a paper, and edges represent the citation relationships between papers.
+The task is to predict the 40 subject areas of arXiv CS papers.
+We collect raw text from [65].
+Wiki-CS is another graph dataset in the OGB collection, specifically designed for node classification tasks within the context of social networks.
+It is based on Wikipedia articles and represents a graph of scientific papers.
+Each node represents a Wikipedia page and each edge represents a reference link.
+Each node’s label corresponds to the category of the entry.
+We collect raw text from [46].
+Ele-Fashion[66] is a node classification dataset derived from the Amazon-Fashion dataset.
+Each node represents a fashion product, and links indicate user co-purchasing behavior.
+Textual features are product titles.
+We collect raw text from[66].
+The task is to predict product categories, simulating recommendation scenarios.
+PCBA is a large-scale chemical bioactivity dataset used for graph machine learning research in cheminformatics and drug discovery.
+Each node represents an atom of a molecule, and the edges represent the chemical bonds between atoms within the molecule.
+The primary tasks in the PCBA dataset are to predict the activity of chemical molecules across various bioactivity assays.
+We collect raw text from [46].
+HIV is a collection of data used to study issues related to the treatment of HIV (Human Immunodeficiency Virus).
+It is commonly used in machine learning and data analysis tasks such as drug discovery, drug-virus interaction prediction, and drug resistance prediction.
+It is a subset of the PubChem BioAssay dataset.
+The task in the HIV dataset involves predicting the activity of chemical molecules against the HIV virus.
+We collect raw text from [46].
+For Cora and PubMed, we measure the performance of all models on the test sets over 10 random splits as suggested in [67] and report the average accuracy.
+For ogbn-arxiv, we follow the official splits [65].
+Following the experimental procedure suggested by OGB, we repeat each experiment for 10 times with random seeds and report the average accuracy.
+For Wiki-CS, we follow the official splits [68] with 20 different training splits and report the average accuracy.
+For PCBA and HIV, we follow the official splits [69], repeat each experiment for 10 times with random seeds and report the average accuracy.
+Table 2: Grid search space.
+Hyper-parameter Search space
+lr {0.01, 0.02, 0.1, 0.2}
+dropout [0.0,0.9]
+weight decay {5e-5, 1e-5, 5e-4, 1e-4, 5e-3}
+$\alpha$ [0.1,0.9]
+$\beta$ [0.1,0.9]
+$\tau$ {0.1,1,10}
+T [0.1,0.9]
+Table 3: Experiment results in supervised learning.
+We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks.
+We highlight the best score on each dataset in bold and the runner-up score with underline.
+Datasets and tasks MLP GAT GIN OFA GraphCL DGI GraphMAE UniGraph BooG
+Node level Cora 80.68 83.57 80.41 77.45 65.79 63.26 78.31 83.02 83.70±0.47
+Pubmed 78.62 80.46 79.26 75.16 72.30 70.81 78.24 81.57 88.51±0.31
+ogbn-arxiv 68.31 70.24 70.55 77.64 60.94 65.77 70.80 73.10 74.57±0.61
+Wiki-CS 65.24 73.61 68.20 78.12 58.44 60.55 64.35 72.85 75.84±0.60
+Ele-Fashion 58.23 62.12 65.24 68.56 55.20 60.17 64.05 70.11 72.34±0.65
+Edge level Cora-link 88.64 90.27 88.85 90.32 81.53 80.86 80.28 90.20 93.11±1.24
+Graph level PCBA 52.31 50.83 60.03 20.89 54.30 55.17 55.67 55.18 58.26±1.27
+HIV 60.15 62.37 70.46 71.47 65.86 62.14 67.04 71.02 74.50±1.42
+Table 4: Experiment results in few-shot learning.
+We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks.
+We highlight the best score on each dataset in bold and the runner-up score with underline.
+Datasets and tasks MLP GAT GIN OFA GraphCL DGI GraphMAE UniGraph BooG
+Node level Cora 1-shot 15.61 20.16 16.24 20.80 33.93 40.45 36.21 39.15 44.47±1.91
+5-shot 47.23 53.21 50.19 60.24 65.28 68.31 68.08 68.55 70.82±1.70
+Pubmed 1-shot 15.92 18.24 19.44 30.67 22.74 20.15 24.76 29.51 33.85±1.78
+5-shot 45.66 50.17 48.79 60.27 35.16 38.61 35.65 58.24 57.56±1.44
+ogbn-arxiv 1-shot 7.24 16.61 20.31 22.38 18.77 16.33 17.68 19.20 20.18±1.74
+5-shot 26.86 36.37 35.65 37.69 35.81 33.45 30.27 42.90 43.31±1.53
+Wiki-CS 1-shot 10.70 21.55 20.33 23.18 20.55 11.25 16.84 19.77 25.11±1.07
+5-shot 35.15 40.86 42.15 47.95 30.34 30.81 34.65 41.12 45.92±0.85
+Ele-Fashion 1-shot 18.50 25.62 26.40 30.22 20.15 22.84 25.95 33.10 35.84±1.21
+5-shot 32.64 42.01 45.32 48.51 33.77 36.44 43.13 52.45 54.92±0.78
+Edge level Cora-link 1-shot 30.06 50.88 52.83 60.27 30.57 46.11 38.82 52.31 58.27±1.36
+5-shot 63.50 76.32 77.47 80.18 55.14 61.28 75.43 79.31 82.41±1.65
+Graph level PCBA 1-shot 7.55 11.04 15.24 18.22 15.11 13.18 16.31 17.02 21.55±1.18
+5-shot 20.16 20.37 28.46 20.14 18.87 21.63 20.62 26.55 28.62±1.54
+HIV 1-shot 30.37 33.67 40.24 45.10 36.16 35.24 33.76 41.12 47.74±1.37
+5-shot 53.84 50.11 57.16 63.25 46.23 50.80 38.57 63.71 65.48±1.02
+Table 5: Experiment results in zero-shot learning.
+We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks.
+We highlight the best score on each dataset in bold and the runner-up score with underline.
+Since OFA is first pre-trained on the citation network Cora with training labels, so we did not report the results of OFA on Cora in zero-shot learning.
+Methods Cora Pubmed ogbn-arxiv Wiki-CS Ele-Fashion Cora-link PCBA HIV
+GraphCL 40.11 30.25 25.05 22.37 26.45 37.86 20.16 28.45
+DGI 38.36 28.46 24.11 23.76 28.16 35.85 21.02 25.11
+GraphMAE 44.93 33.57 23.81 20.44 29.42 45.63 18.67 22.35
+OFA - 35.07 30.86 28.66 33.88 40.80 23.64 50.44
+UniGraph 63.21 66.73 35.88 36.28 30.92 62.61 27.04 48.57
+BooG 69.01±0.41 39.11±0.82 35.24±0.76 36.50±0.84 35.77±1.20 70.24±1.73 25.76±0.84 47.61±1.35
 Baselines.
-We compare BooG with six baselines, which can be categorized into four types.
+We compare BooG with eight baselines, which can be categorized into four types.
 (1) MLP utilizes a multi-layer perceptron to extract deep features individually for each node.
@@
 DGI [37] introduces a self-supervised pre-training task that aims to maximize the mutual information between the local node view and the global graph view.
-(4) Graph foundation models: OFA [14] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs.
+GraphMAE [41] present a masked graph autoencoder that mitigates the issues for generative self-supervised graph learning.
+(4) graph models: OFA [46] utilizes a pre-trained LM to align feature spaces of graphs from different domains, enabling supervised training across all graphs.
+UniGraph We further include recent graph foundation models such as UniGraph [48], which leverages LLMs and GNNs through masked modeling and instruction tuning for generalized graph representations.
+Figure 4: t-SNE visualization on Cora.
+Experimental Setup.
+We implement BooG by PyTorch and conducted the experiments with one A100 GPU.
+The model is optimized by Adam [70].
+We employ TAGs for all methods to unify all input data.
+Specifically, we employ the pre-trained language models (LMs) to encode instance features as input for models.
+BooG and all other self-supervised learning baselines are first pre-trained on the citation network Cora.
+For other self-supervised learning baseline methods, we train a linear classifier on the top of the embeddings derived from the frozen model.
+Then, we evaluate the model performance on the test sets of all seven datasets with different tasks.
+We perform a grid search to tune hyper-parameters based on the validation set.
+Details on the search space of these hyper-parameters can be found in Table 2.
+Figure 5: Ablation study.
 5.2 Performance Results
@@
 For zero-shot learning, different from the above two scenarios, OFA is first pre-trained on the citation network Cora.
-For GraphCL and DGI, predictions are generated by comparing the similarity between the instance representations derived from their models and the class label representations encoded by the LM.
+For GraphCL, DGI and GraphMAE, predictions are generated by comparing the similarity between the instance representations derived from their models and the class label representations encoded by the LM.
 Finally, we evaluate all methods on the test sets of the target datasets.
@@
 However, these methods perform poorly in other domains because they are typically designed for a single graph domain, especially in few-shot learning.
-(3) The two self-supervised learning methods, GraphCL and DGI, exhibit poor domain generalization capabilities in few-shot learning and zero-shot learning, since they only use the raw graph structure.
+(3) The three self-supervised learning methods, GraphCL, DGI and GraphMAE exhibit poor domain generalization capabilities in few-shot learning and zero-shot learning, since they only use the raw graph structure.
 (4) OFA demonstrates a certain degree of domain generalization capabilities in few-shot learning, highlighting the effectiveness of in-context learning.
@@
 Overall, BooG shows strong generalization ability in all scenarios, achieving results comparable to or even better than baseline methods, which proves its effectiveness.
-Table 1: Experiment results in supervised learning.
+Table 6: Experiment results in supervised learning on different pre-training datasets.
 We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks.
-We highlight the best score on each dataset in bold and the runner-up score with underline.
-Datasets and tasks MLP GAT GIN OFA GraphCL DGI BooG
+We highlight the best score on each dataset in bold.
+Pre-training Cora Pubmed ogbn-arxiv Wiki-CS Ele-Fashion Cora-link PCBA HIV
+Cora 83.70±0.47 88.51±0.31 74.57±0.61 75.84±0.60 72.34±0.65 93.11±0.41 58.26±1.27 74.50±1.42
+Pubmed 80.77±0.51 89.36±0.45 74.89±0.46 75.07±0.84 68.02±0.91 90.65±1.37 60.25±1.46 74.11±1.82
+Wiki-CS 76.85±0.13 84.26±0.51 68.42±0.73 79.30±0.37 66.74±1.00 80.16±1.42 58.86±1.33 75.89±1.64
 […]
+5.3 Visualization
+We further evaluate the expressiveness of BooG.
+In particular, we visualize node representations before and after pre-training on Cora with the t-SNE technique [71].
+Colors are employed to differentiate between various node classes, allowing for clear visual distinction.
+Prior to pre-training, the node representations are generated by the language model (LM), which encodes the input data.
+As shown in Figure 4, after undergoing pre-training, the node representations corresponding to different classes become more distinctly separated in the representation space.
+This improvement in separation suggests that the BooG model is able to learn richer and more expressive feature representations, capturing more meaningful distinctions between the node classes.
+Table 7: Experiment results in few-shot learning on different pre-training datasets.
+We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks.
+We highlight the best score on each dataset in bold.
+Pre-training Cora Pubmed ogbn-arxiv Wiki-CS Ele-Fashion Cora-link PCBA HIV
+Cora 1-shot 44.47±1.91 33.85±1.78 20.18±1.74 25.11±1.07 35.84±1.21 58.27±1.36 21.55±1.18 47.74±1.37
+5-shot 70.82±1.70 57.56±1.44 43.31±1.53 45.92±0.85 54.92±0.78 82.41±1.65 28.62±1.54 65.48±1.02
+Pubmed 1-shot 38.91±1.26 40.57±1.31 24.33±1.52 24.86±1.44 32.16±1.10 54.35±1.45 23.82±1.45 48.83±1.44
+5-shot 68.25±1.04 65.86±1.27 45.16±1.31 46.03±1.15 49.03±0.92 79.30±1.58 35.41±1.61 65.80±1.32
+Wiki-CS 1-shot 31.42±1.34 28.46±1.21 21.65±1.82 31.62±1.03 29.70±1.05 49.33±1.42 23.67±1.48 47.16±1.35
+5-shot 54.28±1.77 53.81±1.45 43.88±1.64 62.16±1.15 50.80±0.85 60.36±1.85 36.12±1.53 64.93±1.48
 […]
+5-shot 60.16±1.84 58.32±1.41 48.26±1.15 60.85±1.37 52.45±0.94 80.30±1.14 39.36±1.28 68.30±1.24
+Table 8: Experiment results in zero-shot learning on different pre-training datasets.
+We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks.
+We highlight the best score on each dataset in bold.
+Pre-training Cora Pubmed ogbn-arxiv Wiki-CS Ele-Fashion Cora-link PCBA HIV
+Cora 69.01±0.41 39.11±0.82 35.24±0.76 36.50±0.84 35.77 ±1.20 70.24±1.73 25.76±0.84 47.61±1.35
+Pubmed 63.21±1.14 66.73±1.27 35.88±1.43 36.28±0.79 30.92±0.93 62.61±1.85 27.04±0.78 48.57±1.65
+Wiki-CS 43.81±0.82 45.64±1.12 30.17±1.52 57.32±0.67 29.42±0.86 60.21±1.57 27.13±0.85 47.62±1.37
 […]
-[…]
-[…]
-[…]
-[…]
-Table 2: Experiment results in few-shot learning.
-We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks.
-We highlight the best score on each dataset in bold and the runner-up score with underline.
-Datasets and tasks MLP GAT GIN OFA GraphCL DGI BooG
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 3: Experiment results in zero-shot learning.
-We report accuracy (%) for node/edge tasks and ROC-AUC score (%) for graph tasks.
-We highlight the best score on each dataset in bold and the runner-up score with underline.
-Since OFA is first pre-trained on the citation network Cora with training labels, so we did not report the results of OFA on Cora in zero-shot learning.
-Datasets and tasks Cora Pubmed ogbn-arxiv Wiki-CS Cora-link PCBA HIV
-[…]
-DGI 38.36 28.46 24.11 23.76 35.85 21.02 25.11
-[…]
-[…]
-Figure 4: t-SNE visualization on Cora.
-Figure 5: Ablation study.
-5.3 Visualization
-We further evaluate the expressive power of BooG.
-We visualize node representations before and after pre-training on Cora with the t-SNE technique [24].
-Colors are used to distinguish different node classes.
-Before pre-training, the node representations are encoded by the LM.
-As shown in Figure 5, the node representations of different classes are more clearly separated after pre-training.
-This indicates that BooG can learn more expressive representations.
 5.4 Ablation Study
-The ablation study is done to understand the importance of main components of BooG in supervised learning and 5-shot learning scenarios on Cora and HIV.
+The ablation study is performed to understand the importance of main components of BooG in supervised learning and 5-shot learning scenarios.
+We select Cora and HIV as two representative datasets.
 The model is pre-trained on Cora.
@@
 Further, the decline in performance exhibited by both BooG-nc and BooG-na highlights the importance of super nodes.
-In other words, the super nodes significantly enhance model generalizability, especially in graph domain transfer.
+In other words, the super nodes significantly enhance model generalizability, especially in graph domain transfer, further demonstrated the effectiveness of BooG.
+5.5 Performance on different pre-training datasets
+We next evaluate the model generalizability across various learning scenarios and downstream tasks based on different pre-training datasets.
+BooG is first pre-trained on Cora/Pubmed/Wiki-CS and their union.
+In the supervised learning setting, the parameters of pre-trained models are frozen and the MLP is trained in a supervised manner.
+In the few-shot learning setting, we freeze the parameters of the pre-trained methods and train their MLPs using a $N$-way $K$-shot task.
+We evaluate BooG on the test sets of the target datasets.
+Table 6, Table 7 and Table 8 report the results in supervised learning, few-shot learning, and zero-shot learning, respectively.
+From these tables, we observe that: (1) BooG achieves good performance on various pre-training datasets.
+(2) Compared to other target datasets, BooG demonstrates better generalization on the pre-training source dataset.
+(3) Joint pre-training on multiple datasets performs better on unseen-domain datasets compared to pre-training on a single dataset.
+Overall, BooG shows strong generalization ability in all scenarios on different pre-training datasets, which proves its effectiveness.
+Figure 6: Hyper-parameter sensitivity analysis.
+Table 9: Comparison of GPU hours and performance on Cora. All time measured on a single A100 GPU.
+Method Pre-training Downstream Training Inference
+GAT - 23 min (supervised) / 5 min (5-shot) 1 min
+GraphCL 52 min 18 min (supervised) / 4 min (5-shot) 1 min
+GraphMAE 63 min 25 min (supervised) / 5 min (5-shot) 1 min
+UniGraph 2.8 h 8 min (supervised) / 5 min (5-shot) 2 min
+BooG 43 min 5 min (supervised) / 2 min (5-shot) 1 min
+5.6 Sensitivity Analysis
+In this section, we analyze the sensitivity of BooG to the hyper-parameters $\alpha$ and $\beta$, which controls the influence of the class label embedding in constructing super nodes.
+Specifically, $\alpha$ determines the balance between the anchor node’s original representation and the class label information when forming the super node embedding.
+A larger $\alpha$ places more emphasis on the class label prior, while a smaller $\alpha$ makes the super node more similar to the original anchor node embedding.
+As shown in Figure 6, the performance of BooG exhibits sensitivity to the choice of $\alpha$.
+Moderate values of $\alpha$ typically yield better results, as they allow the model to effectively balance the anchor node’s instance-specific semantics and the global class-level prior.
+If $\alpha$ is too small, the super node degenerates to the original node representation, reducing the benefit of class-conditioned modeling.
+Conversely, if $\alpha$ is too large, the model overemphasizes class priors and may ignore useful instance-specific signals.
+Unlike vanilla GNNs, BooG explicitly leverages class priors via super nodes, and $\alpha$ controls this mechanism.
+For hyper-parameter $\beta$, BooG consistently performs well across a wide range of values, which demonstrates that $\beta$ does not significantly impact the performance of BooG.
+5.7 Efficiency Analysis
+To further address practical efficiency, we compare the GPU time required for both pre-training and downstream tasks on the Cora dataset.
+Table 9 reports the results.
+We observe that BooG requires significantly less pre-training time (only 43 minutes) compared to existing graph pre-training methods such as GraphCL (52 minutes), GraphMAE (63 minutes), and UniGraph (2.8 hours).
+This efficiency is attributed to BooG’s lightweight design, which avoids costly full-graph augmentation or reconstruction processes and instead leverages text semantics and localized graph neighborhoods.
+For downstream tasks, BooG also achieves fast tuning and inference.
+The MLP-based classifier on BooG’s representations requires only 5 minutes for supervised tuning, and even faster under 5-shot settings (2 minute).
+The inference process relies on simple similarity matching, making BooG extremely efficient at deployment.
+These results highlight that BooG provides a favorable balance between efficiency and generalization.
 6 Conclusion
-In this paper, we boost graph foundation model from structural perspective and propose BooG.
-The model constructs virtual super nodes to unify structural characteristics of graph data from different domains.
-Specifically, the virtual super nodes fuse the information of both anchor nodes and class labels, where each anchor node captures the information of a node instance or a graph instance to be classified.
-Instead of utilizing the raw graph structure, we introduce super nodes and connect them to all the nodes within the neighborhood of an anchor by virtual edges.
-This new linkage allows for effective information aggregation while disregarding the difference in structural connectivity across domains.
-In this way, cross-domain model generalization can be boosted from structural perspective.
-Additionally, we propose a novel pre-training objective based on contrastive learning, which learns more expressive representations for graph data and generalizes effectively to different domains and downstream tasks.
-Our extensive experiments and ablation studies validate the effectiveness of BooG.
-Appendix A Datasets
-In this section, we introduce the datasets used in the experiments. The statistics of datasets are listed in Table 4, and we describe each dataset in detail below:
-Cora is a homophilous graphs commonly used as benchmarks.
-It consists of 2708 scientific publications classified into one of 7 classes.
-In this dataset, each node represents a scientific paper and each edge represents a citation.
-We collect raw text from [5].
-Pubmed is a homophilous graphs commonly used as benchmarks.
-It consists of 19,717 scientific publications from PubMed database pertaining to diabetes classified into one of 3 classes.
-We collect raw text from [5].
-ogbn-arXiv is an academic citation network graph dataset, consisting of papers from arXiv and their citation relationships.
-Each node represents a paper, and edges represent the citation relationships between papers.
-The task is to predict the 40 subject areas of arXiv CS papers.
-We collect raw text from [8].
-Wiki-CS is an Internet link network where each node represents a Wikipedia page and each edge represents a reference link.
-Each node’s label corresponds to the category of the entry.
-We collect raw text from [14].
-PCBA is a large-scale chemical bioactivity dataset used for graph machine learning research in cheminformatics and drug discovery.
-Each node represents an atom of a molecule, and the edges represent the chemical bonds between atoms within the molecule.
-The primary tasks in the PCBA dataset are to predict the activity of chemical molecules across various bioactivity assays.
-We collect raw text from [14].
-HIV is a subset of the PubChem BioAssay dataset, focusing on predicting the activity of chemical molecules against the HIV virus.
-The task in the HIV dataset involves predicting the activity of chemical molecules against the HIV virus.
-We collect raw text from [14].
-For Cora and PubMed, we measure the performance of all models on the test sets over 10 random splits as suggested in [20] and report the average accuracy.
-For ogbn-arxiv, we follow the official splits [8].
-Following the experimental procedure suggested by OGB, we repeat each experiment for 10 times with random seeds and report the average accuracy.
-For Wiki-CS, we follow the official splits [19] with 20 different training splits and report the average accuracy.
-For PCBA and HIV, we follow the official splits [32], repeat each experiment for 10 times with random seeds and report the average accuracy.
-The statistics and details of these datasets can be found in Table 4.
-Table 4: Statistics of all text-attributed graph datasets.
-Dataset Domain Task #Graphs Avg.#Nodes Avg.#Edges Raw Texts
-Cora Citation Node/Link 1 2,708 10,556 paper titles and abstracts
-Pubmed Citation Node/Link 1 19,717 44,338 paper titles and abstracts
-ogbn-arxiv Citation Node 1 169,343 1,166,243 paper titles and abstracts
-Wiki-CS Web link Node 1 11,701 216,123 wikipedia entry names and contents
-PCBA Molecule Graph 437,929 26.0 28.1 textual descriptions of atoms/bonds
-HIV Molecule Graph 41,127 25.5 27.5 textual descriptions of atoms/bonds
-Appendix B Experimental Setup
-We implement BooG by PyTorch and conducted the experiments with one A100 GPU.
-The model is optimized by Adam [10].
-We employ TAGs for all methods to unify all input data.
-Specifically, we employ the pre-trained language models (LMs) to encode instance features as input for models.
-BooG and all other self-supervised learning baselines are first pre-trained on the citation network Cora.
-For other self-supervised learning baseline methods, we train a linear classifier on the top of the embeddings derived from the frozen model.
-Then, we evaluate the model performance on the test sets of all six datasets with different tasks.
-We perform a grid search to tune hyper-parameters based on the validation set.
-Details on the search space of these hyper-parameters can be found in Table 5.
-Table 5: Grid search space.
-Hyper-parameter Search space
-lr {0.01, 0.02, 0.1, 0.2}
-dropout [0.0,0.9]
-weight decay {5e-5, 1e-5, 5e-4, 1e-4, 5e-3}
-$\alpha$ [0.1,0.9]
-$\beta$ [0.1,0.9]
-$\tau$0.1,1,10}
-T [0.1,0.9]
-Appendix C Sensitivity Analysis
-Here we study the sensitivity of two key hyper-parameters: $\alpha$ and $\beta$.
-The results are shown in Figure 6.
-From the figure, we see that the hyper-parameter $\alpha$, which represents the importance of class labels when constructing super nodes, has a significant impact on the model’s performance.
-If the weight of the class labels is too small, the class labels will not be effective, leading to a decline in model performance.
-Conversely, if the weight of the class labels is too large, the representations of the instances may lose their individuality, also resulting in a performance drop.
-For hyper-parameter $\beta$, BooG consistently performs well across a wide range of values, which demonstrates that $\beta$ does not significantly impact the performance of BooG.
-Figure 6: Hyper-parameter sensitivity analysis.
-Appendix D Broader Impacts
-The primary objective of this study is to chieves robust performance and generalization capabilities across various graph domains and tasks with a graph foundation model.
-The results indicate that the proposed model contribute to enhance the model generalizability, which can have significant implications in various fields such as social network analysis and molecular property prediction.
-However, this work may also have some negative impacts.
-As deep neural network models become more widespread, they can be prone to misuse, potentially resulting in detrimental outcomes.
-Additionally, the privacy of training data remains an issue, as sensitive information may be unintentionally leaked through the model.
-We could incorporate responsible AI techniques for GNNs into our model to avoid the issues of misuse and privacy.
-Appendix E Limitations
-We discuss the limitations and future directions of our work in this section.
-This method is currently only suitable for graph classification, node classification and link prediction tasks.
-In future work, we aim to extend our method to a wider range of graph tasks.
-In addition, our method does not take advantage of the textual attributes of the edges, and we also aim to extend our method to more data attributes.
+In this paper, we boost graph models from structural perspective and propose BooG.
+BooG boosts model’s cross-domain and cross-task generalization ability from two key aspects.
+First, BooG unifies graph attributes by encoding node texts of TAGs and class label descriptions into a shared space using a pre-trained LM, thereby equipping the model to handle graph data from diverse domains.
+More critically, BooG unifies graph structures by introducing the concept of sub-graphs and super nodes.
+Sub-graphs offer a consistent task template across all levels of tasks, consisting of anchor nodes and their neighbor nodes.
+The super nodes, along with virtual edges, establish a standardized aggregation mechanism that fuses rich information from neighborhoods and associated class labels, accommodating graph structural characteristics inherent to different domains.
+These designs integrate a supernode-centered graph structure that encapsulates label preferences, making BooG compatible with different graph domains and task levels.
+Our extensive experiments on seven datasets, covering a range of domains and tasks, show that BooG outperforms state-of-the-art competitors in most cases.
+{acknowledgement}
+This work is supported by National Natural Science Foundation of China No. 62202172.